OpenStudy (anonymous):

Integration Question, w/ Trig substitution. My question is, can I use trig sub. with a cubed root? I.e \[\sqrt[3]{x^2+1} \]

OpenStudy (anonymous):

I've got all the work done. I guess I'm just verifying that it is indeed alright. For instance in this case to use \[x=\tan \theta\]

OpenStudy (anonymous):

\(cos^2\theta + sin^2\theta = 1\), divide by \(cos^2\theta\) \(1 + tan^2\theta = sec^2\theta\), now if u call: \(x = tan\theta\), we get: \(\sqrt[3]{sec^2\theta}\)

OpenStudy (anonymous):

and \(dx = sec^2\theta d\theta\)

OpenStudy (anonymous):

is this answer your question?

OpenStudy (anonymous):

I guess my question was a "non-question" now that I read it again. Trig sub is just ideal for situations such as square roots because you can easily get squared values through trig identities. Sadly I made a mistake in my work where, out of habit of using the square root for trig sub, I took the root of squared tan(theta) which is incorrect considering it's a cubed root function. Thanks for making me see it. Cheers

OpenStudy (anonymous):

Np ;)